# Universal projection theorems with applications to multifractal analysis and the dimension of every ergodic measure on self-conformal sets simultaneously

### v2

@@ 1. Introduction
(1.3) $\gamma\left(\left\{V\in\operatorname{Gr}(d,n):\text{ there exists }\mu\in{\mathcal{M}}\text{ such that}\dim_{H}(P_{V}\mu)<\min\left\{d,\dim_{H}\mu\right\}\right\}\right)=0.$
Clearly, this question is relevant only for uncountable families ${\mathcal{M}}$ and it is easy to find examples of collections ${\mathcal{M}}$ for which (1.3) does not hold (see Example 3.2). Our main result (Theorem 2.7) implies that (1.3) holds if ${\mathcal{M}}$ satisfies a condition which we call relative dimension separability - see Definition 2.3. We can verify it in two important cases of dynamical origin, proving that
@@
    (1.3) holds if ${\mathcal{M}}$ is the family of all ergodic invariant measures on a self-conformal set satisfying the strong separation condition - see Theorem 3.5,
@@
    (1.3) holds if ${\mathcal{M}}$ is the family of all Gibbs measures (corresponding to all Hölder continuous potentials) on a self-conformal set (without any separation conditions) - see Theorem 3.6.
For parametrized families of $C^{1+\theta}$ conformal IFS ${\mathcal{F}}_{\lambda}$ satisfying the transversality condition, we obtain a version of (1.4) which holds simultaneously for all ergodic measures on ${\mathcal{A}}^{\mathbb{N}}$:
(1.4) $\eta\left(\left\{\lambda\in U:\text{ there exists }\mu\in{\mathcal{E}}_{\sigma}({\mathcal{A}}^{\mathbb{N}})\text{ such that}\dim_{H}(\Pi_{\lambda}\mu)<\min\left\{d,\frac{h(\mu)}{\chi(\mu,\lambda)}\right\}\right\}\right)=0,$
Our motivation for studying simultaneous projection result like (1.4) is twofold. First and foremost, this approach has novel applications to multifractal analysis. We verify that the multifractal formalism holds on the full spectrum interval for typical self-similar measures with overlaps for almost every translation parameter - see Theorems 3.14 and 8.3. In dimension one, we prove that given a self-similar system $\{x\mapsto\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ on the line with similarity dimension $s_{0}$ smaller than one, then Lebesgue almost every choice of translations $t_{i}$ has the property that the multifractal formalism holds on the full spectrum interval $\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$ for every self-similar measure corresponding to a probability vector $p=(p_{i})_{i\in{\mathcal{A}}}$ simultaneously. According to our best knowledge, it is the first result on the multifractal spectrum of self-similar measures without separation conditions which covers the full spectrum and improves in this regard recent results of Barral and Feng [BF21]. See Sections 3.4 and 8 for details. It is easy to extend our results to the Birkhoff spectra of continuous potentials. Our proofs of the multifractal results are based on a study of a separation property called the exponential distance from the enemy (EDE), which in terms of the natural projection $\Pi_{\lambda}$ can be seen as a nearly bi-Lipschitz (pointwise $\alpha$-Hölder for every $\alpha\in(0,1)$) continuity property of the inverse to $\Pi_{\lambda}$ at almost every point - see Theorem 3.11 and Proposition 3.10. The latter property is obtained in a general version which can be applied also to orthogonal projections in terms of the Assouad dimension, giving a simultaneous version of [BGŚ23, Theorem 1.7.(iii)].
The second goal of our study is to bridge a gap between results obtained with the use of transversality technique and more recent approach using additive combinatorics. In the case when the IFS consists of similarity mappings and the measure $\mu$ is Bernoulli (i.e. $\Pi\mu$ is a self-similar measure), the breakthrough result of Hochman [Hoc14] significantly strengthened results like (1.2). Namely, Hochman showed that if the IFS $\Phi=\{f_{i}(x)=\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ satisfies the exponential separation condition (ESC) then (1.2) holds for every Bernoulli measure $\mu$. Furthermore, under sufficiently smooth (analytic) and non-degenerate parametrization, the ESC holds for almost every parameter (and often outside of a set of parameters of non-full dimension). Later, Jordan and Rapaport [JR21], relying on the result of Shmerkin [Shm19] on $L^{q}$ dimension of self-similar measures, showed that equality in (1.2) holds for every ergodic shift-invariant measure simultaneously provided that the system satisfies ESC. Our result (1.4) shows that one can achieve the same simultaneity among ergodic measures using the transversal technique. While transversality remains inferior to additive combinatorics methods in some regards (e.g. transversality is rarely able to cover the whole parameter range), it is able to produce similar results like (1.2) for families of general non-linear $C^{1+\theta}$ IFS. As for orthogonal projections, Bruce and Jin showed in [BJ19, Theorem 1.3] that given a self-conformal IFS on $\mathbb{R}^{n}$ which satisfies certain irrationality condition for the derivatives, one has $\dim_{H}P_{V}\mu=\min\{d,\dim_{H}\mu\}$ for every $V\in\operatorname{Gr}(d,n)$ and every Gibbs measure $\mu$. This extends results of Hochman and Shmerkin [HS12]. Our Theorem 3.6 does not require any irrationality condition at the cost of allowing the dimension formula to fail for a zero-measure set of projections.
We prove our main result, which we call a universal projection theorem in a general setting of a family of projection maps which can encompass both orthogonal projections and natural projections corresponding to a family of iterated projection schemes, as well as some other families like non-autonomous IFS, where in each iterate we choose different IFS with the same parametrization, see Nakajima [Nak24] or when in each iterate we add an independent, identically distributed error, see Jordan, Pollicott and Simon [JPS07], Koivusalo [Koi14] and Liu and Wu [LW03]. This setting is also capable of covering certain suitable parametrizations of delay-coordinate maps (at least for aperiodic systems [Rob11, Chapter 14]), which are used for obtaining time-delayed embeddings of dynamical systems [SYC91, Rob11], studied also in the probabilistic context (see e.g. [SY97, BGŚ20, BGŚ24]) but we do not pursue this direction in the present work. Our general setting is similar to, but slightly different than, some other generalized projection schemes that appeared in the literature, e.g. in [Sol98, PS00] or [BSS23, Section 6.6]. Finally, the name universal projection theorem is inspired by seminal results and constructions known in information theory as universal source codings, which give compression algorithms operating in an optimal rate for any source distribution in a given class, without any prior knowledge of this distribution, see e.g. [CT06, Chapter 13]. A famous example is the Lempel-Ziv coding [ZL77, ZL78] achieving optimal rate for any stationary ergodic source distribution [CT06, Theorem 13.5.3].
@@ 2. Main result: a universal projection theorem
We shall now describe the setting in which we will prove a universal projection theorem. Let $X$ be a compact topological space (the phase space) and let $U$ be a hereditary Lindelöf topological space111topological space $U$ is hereditary Lindelöf if every subset has the property that its every open cover has a countable subcover. Every separable metric space is hereditary Lindelöf, see e.g. [Eng89]. (the parameter space). Let $\eta$ be a locally finite Borel measure on $U$. For $\lambda\in U$ let $\rho_{\lambda}$ be a metric on $X$ compatible with its topology and let $B_{\lambda}(x,r)$ denote the open $r$-ball in metric $\rho_{\lambda}$. Fix $d\in{\mathbb{N}}$ and for each $\lambda\in U$ consider a map $\Pi_{\lambda}:X\to\mathbb{R}^{d}$ (projection). Let ${\mathcal{M}}_{\mathrm{fin}}(X)$ denote the set of all finite Borel measures on $X$ and consider a collection ${\mathcal{M}}\subset{\mathcal{M}}_{\mathrm{fin}}(X)$. Our goal is to study projections $\Pi_{\lambda}\mu,\lambda\in U,\mu\in{\mathcal{M}}$ and provide projection theorems which hold for $\eta$-a.e. $\lambda\in U$ and all measures $\mu\in{\mathcal{M}}$ simultaneously. A crucial property needed for the proofs is a certain separability-like condition for the set ${\mathcal{M}}$ and a transversality condition for the family $\Pi_{\lambda}$. For the former one, we make the following definitions.
@@ 3.2. Conformal IFS
Let ${\mathcal{A}}$ be a finite set and let $V\subset\mathbb{R}^{n}$ be a compact connected set with $V=\overline{\operatorname{Int}(V)}$. For each $i\in{\mathcal{A}}$, let $f_{i}:V\to V$ be a conformal $C^{1+\theta}$ map such that $10<\|f^{\prime}_{i}(x)\|<$ 1 for every $x\in V$. We call the collection ${\mathcal{F}}=(f_{i})_{i\in{\mathcal{A}}}$ a conformal $C^{1+\theta}$ IFS. Let $\Sigma={\mathcal{A}}^{\mathbb{N}}$ be the symbolic space over the alphabet ${\mathcal{A}}$. We can associate to ${\mathcal{F}}$ a natural projection map $\Pi_{\mathcal{F}}:\Sigma\to V$ defined as
@@
[…]
$\dim_{H}(\mu,\rho_{\mathcal{F}})=\frac{h(\mu)}{\chi(\mu,{\mathcal{F}})},$
@@
$\dim_{H}P_{V}\mu=\min\{d,\dim_{H}\mu\}\text{ simultaneously for all ergodic measures }\mu\text{ on }\Lambda.$
@@
$\dim_{H}P_{V}\mu=\min\{d,\dim_{H}\mu\}\text{ simultaneously for all Gibbs measures }\mu\text{ on }\Lambda.$
@@
Let ${\mathcal{A}}$ be a finite set and let $V\subset\mathbb{R}^{d}$ be a compact connected set with $V=\overline{\operatorname{Int}(V)}$. For each $\lambda\in U$, where $U$ is a (hereditary Lindelöf) topological space, let ${\mathcal{F}}^{\lambda}=(f_{i}^{\lambda})_{i\in{\mathcal{A}}}$ be a conformal $C^{1+\theta}$ IFS on $V$. Assume that there exist $10<\gamma_{1}<\gamma_{2}<$ 1 such that $\gamma_{1}\leq\|(f^{\lambda}_{i})^{\prime}(x)\|\leq\gamma_{2}$ for every $x\in V,\lambda\in U$. Moreover, assume that the map $U\ni\lambda\mapsto{\mathcal{F}}_{\lambda}$ is continuous, where the distance between ${\mathcal{F}}^{\lambda_{1}}$ and ${\mathcal{F}}^{\lambda_{2}}$ is defined as $\max\limits_{i\in{\mathcal{A}}}\left(\|f^{\lambda_{1}}_{i}-f^{\lambda_{2}}_{i}\|+\|(f^{\lambda_{1}}_{i})^{\prime}-(f^{\lambda_{2}}_{i})^{\prime}\|+\sup\limits_{x\neq y\in V}\frac{\|(f^{\lambda_{1}}_{i})^{\prime}(x)-(f^{\lambda_{2}}_{i})^{\prime}(y)\|}{|x-y|^{\theta}}\right)$. We call a parametrized family ${\mathcal{F}}^{\lambda},\lambda\in U$ a continuous family of $C^{1+\theta}$ conformal IFS. Let $\Sigma={\mathcal{A}}^{\mathbb{N}}$ be the symbolic space, and for each $\lambda\in U$, let $\Pi_{\lambda}:=\Pi_{{\mathcal{F}}^{\lambda}}:\Sigma\to\mathbb{R}^{d}$ be the corresponding natural projection map (3.2). For each $\lambda\in U$, let $\rho_{\lambda}:=\rho_{{\mathcal{F}}^{\lambda}}$ be the metric on $\Sigma$ corresponding to ${\mathcal{F}}^{\lambda}$, as defined in (3.3). Under these assumptions families $\{\Pi_{\lambda}:\lambda\in U\}$ and $\{\rho_{\lambda}:\lambda\in U\}$ satisfy assumptions (A1) and (A2). Property (A1) follows by the bounded distortion property and the distortion continuity, see for example Nakajima [Nak24, Section 3] in the more general non-autonomous case, while (A2) follows directly from the definition of metric $\rho_{\lambda}$.
@@
Let ${\mathcal{F}}=(f_{i})_{i\in{\mathcal{A}}}$ be a conformal $C^{1+\theta}$ IFS on a compact connected set $V\subset\mathbb{R}^{d}$ as described in Example 3.3 and assume that $\max\limits_{i\neq j\in{\mathcal{A}}}\|f_{i}^{\prime}\|+\|f_{j}^{\prime}\|<1$. Let $U=\{(\lambda_{i})_{i\in{\mathcal{A}}}\in(\mathbb{R}^{d})^{{\mathcal{A}}}:f_{i}(V)+\lambda_{i}\subset V\}$, and let $\eta$ be the normalized Lebesgue measure on $U$. Then ${\mathcal{F}}^{\lambda}=\{f_{i}^{\lambda}=f_{i}+\lambda_{i}\}_{i\in{\mathcal{A}}},\lambda\in U$ is a continuous family of $C^{1+\theta}$ conformal IFS. A family of natural projections $\{\Pi_{\lambda}:\lambda\in U\}$ and corresponding metrics $\{\rho_{\lambda}:\lambda\in U\}$ as in Example 3.7 satisfy (A1) - (A3) with measure $\eta$. See for example [BSS23, Theorem 14.5.2] for the proof in the case $d=1$, which extends in a straightforward manner to higher dimensions. ∎
Before formulating our main result on parametrized families of IFS satisfying the transversality condition, let us interpret the nearly bi-Lipschitz condition for natural projection maps as a separation condition for cylinders. Given an infinite word $\omega=(\omega_{1},\omega_{2},\ldots)\in\Sigma$, let $\omega|_{n}=(\omega_{1},\ldots,\omega_{n})$ denote its restriction to the first $n$ coordinates, while given a finite word $\omega\in\Sigma^{*}$, let $|\omega|$ denote its length and let $[\omega]=\{\tau\in\Sigma:\tau|_{n}=\omega\}$ be the corresponding cylinder.
@@
Therefore, for a measure $\mu\in{\mathcal{M}}_{\mathrm{fin}}(\Sigma)$, $\mu$-a.e. $\omega\in\Sigma$ satisfies EDE if and only if $\Pi_{\mathcal{F}}$ is $\mu$-nearly bi-Lipschitz in metric $\rho_{\mathcal{F}}$. For the proof of Proposition 3.10 see Section 7.3. Now we can state our main result on transversal families of conformal IFS.
@@
[…]

### v3

@@ 1.4. Applications II - dimension of every ergodic measure on self-conformal sets simultaneously
(1.7) $\eta\left(\left\{\lambda\in U:\text{ there exists }\mu\in{\mathcal{E}}_{\sigma}({\mathcal{A}}^{\mathbb{N}})\text{ such that}\dim_{H}(\Pi_{\lambda}\mu)\neq\min\left\{d,\frac{h(\mu)}{\chi(\mu,\lambda)}\right\}\right\}\right)=0,$
The setting of our general result is wide enough to encompass also other cases, like non-autonomous IFS, where in each iterate we choose different IFS with the same parametrization (see Nakajima [Nak24]) or when in each iterate we add an independent, identically distributed error (see Jordan, Pollicott and Simon [JPS07], Koivusalo [Koi14] and Liu and Wu [LW03]). More details are given in Section 3.3.
A strong dimension result holding simultaneously for all ergodic measures on self-similar sets in $\mathbb{R}$ was obtained by Jordan and Rapaport [JR21, Theorem 1.1]. They showed that if an IFS consisting of similarities on $\mathbb{R}$ satisfies the Exponential Separation Condition (ESC), then (1.6) holds for all ergodic measures $\mu$ on $\mathcal{A}$$^{\mathbb{N}}$. Our result (1.7) extends this to typical non-linear IFS on $\mathbb{R}^{n}$ satisfying the transversality condition. Very recently, Rapaport [Rap25] extended considerably Hochman’s result [Hoc14], proving that (1.6) holds for any IFS on $\mathbb{R}$ consisting of analytic contractions satisfying ESC and every Bernoulli measure $\mu$ on $\mathcal{A}$$^{\mathbb{N}}$. It however seems a challenging problem to extend Rapaport’s result to all ergodic measures in a fashion similar to [JR21], as the latter relies crucially on Shmerkin’s result on $L^{q}$ dimension of self-similar measures with ESC [Shm19], which does not hold for IFS consisting of analytic maps, see the discussion in [BKT25, Section 1.2.2]. Our result (1.7) is able to deal with all ergodic measures and requires weaker regularity ($C^{1+\theta}$ instead of analyticity) than Rapaport’s [Rap25], at the cost of assuming transversality instead of ESC.
1.5. Applications III - orthogonal projections of ergodic measures on self-conformal sets
Applications of the universal projection theorem (Theorem 2.7) require verifying the relative dimension separability property of a given family $\mathcal{M}$} of measures on a set $X$. In particular, we can verify it in two important cases of dynamical origin, related to the IFS theory discussed above:
@@ 1.5. Applications III - orthogonal projections of ergodic measures on self-conformal sets
    for $\mathcal{M}$} being the family of all ergodic invariant measures on a self-conformal set $X$ satisfying the strong separation condition,
@@
    for $\mathcal{M}$} being the family of all Gibbs measures (corresponding to all Hölder continuous potentials) on a self-conformal set $X$ (without any separation conditions).
This allows us to conclude that Marstand’s projection theorem (1.1) holds simultaneously for $\mu\in{\mathcal{M}}$, with $\mathcal{M}$} as in i and ii. See Theorems 3.5 and 3.6 for precise statements. This can be compared with a result of Bruce and Jin [BJ19, Theorem 1.3], showing that given a self-conformal IFS on $\mathbb{R}^{n}$ which satisfies certain irrationality condition for the derivatives, one has $\dim_{H}P_{V}\mu=\min\{d,\dim_{H}\mu\}$ for every $V\in\operatorname{Gr}(d,n)$ and every Gibbs measure $\mu$ (extending the results of Hochman and Shmerkin [HS12]). Our Theorem 3.6 does not require any irrationality condition, at the cost of allowing the dimension formula to fail for a zero-measure set of projections.
@@ 2. Main result: a universal projection theorem
We shall now describe the setting in which we will prove a universal projection theorem. Let $X$ be a compact topological space (the phase space) and let $U$ be a hereditary Lindelöf topological space333topological space $U$ is hereditary Lindelöf if every subset has the property that its every open cover has a countable subcover. Every separable metric space is hereditary Lindelöf, see e.g. [Eng89]. (the parameter space). Let $\eta$ be a locally finite Borel measure on $U$. For $\lambda\in U$ let $\rho_{\lambda}$ be a metric on $X$ compatible with its topology and let $B_{\lambda}(x,r)$ denote the open $r$-ball in metric $\rho_{\lambda}$. Fix $d\in{\mathbb{N}}$ and for each $\lambda\in U$ consider a map $\Pi_{\lambda}:X\to\mathbb{R}^{d}$ (projection). Let ${\mathcal{M}}_{\mathrm{fin}}(X)$ denote the set of all finite Borel measures on $X$ and consider a collection ${\mathcal{M}}\subset{\mathcal{M}}_{\mathrm{fin}}(X)$. Our goal is to study projections $\Pi_{\lambda}\mu,\lambda\in U,\mu\in{\mathcal{M}}$ and provide projection theorems which hold for $\eta$-a.e. $\lambda\in U$ and all measures $\mu\in{\mathcal{M}}$ simultaneously. A crucial property needed for the proofs is a certain separability-like condition for the set $\mathcal{M}$} and a transversality condition for the family $\Pi_{\lambda}$. For the former one, we make the following definitions.
@@ 3.2. Conformal IFS
Let $\mathcal{A}$} be a finite set and let $V\subset\mathbb{R}^{n}$ be a compact connected set with $V=\overline{\operatorname{Int}(V)}$. For each $i\in{\mathcal{A}}$, let $f_{i}:V\to V$ be a conformal $C^{1+\theta}$ map such that $0<\|f^{\prime}_{i}(x)\|<1$ for every $x\in V$. We call the collection ${\mathcal{F}}=(f_{i})_{i\in{\mathcal{A}}}$ a conformal $C^{1+\theta}$ IFS. Let $\Sigma={\mathcal{A}}^{\mathbb{N}}$ be the symbolic space over the alphabet $\mathcal{A}$}. We can associate to $\mathcal{F}$} a natural projection map $\Pi_{\mathcal{F}}:\Sigma\to V$ defined as
@@
where $\|\cdot\|$ is the supremum norm on $V$. It is easy to check that $\rho_{\mathcal{F}}$ is a metric on $\Sigma$ and the natural projection map $\Pi_{\mathcal{F}}:\Sigma\to\mathbb{R}^{n}$ is Lipschitz in $\rho_{\mathcal{F}}$. If $\mathcal{F}$} satisfies the Strong Separation Condition, then $\Pi_{\mathcal{F}}$ is bi-Lipschitz in $\rho_{\mathcal{F}}$ (see [BSS23, Section 14.2]). The metric $\rho_{\mathcal{F}}$ is also significant as natural dynamical invariants of the system can be expressed in terms of dimensions calculated with respect to $\rho_{\mathcal{F}}$. Namely, an ergodic shift-invariant measure $\mu$ on $\Sigma$ is exact-dimensional with respect to $\rho_{\mathcal{F}}$ and satisfies
(3.4) $\dim_{H}(\mu,\rho_{\mathcal{F}})=\frac{h(\mu)}{\chi(\mu,{\mathcal{F}})},$
@@
[…]
@@
[…]
@@
Let $\mathcal{A}$} be a finite set and let $V\subset\mathbb{R}^{d}$ be a compact connected set with $V=\overline{\operatorname{Int}(V)}$. For each $\lambda\in U$, where $U$ is a (hereditary Lindelöf) topological space, let ${\mathcal{F}}^{\lambda}=(f_{i}^{\lambda})_{i\in{\mathcal{A}}}$ be a conformal $C^{1+\theta}$ IFS on $V$. Assume that there exist $0<\gamma_{1}<\gamma_{2}<1$ such that $\gamma_{1}\leq\|(f^{\lambda}_{i})^{\prime}(x)\|\leq\gamma_{2}$ for every $x\in V,\lambda\in U$. Moreover, assume that the map $U\ni\lambda\mapsto{\mathcal{F}}_{\lambda}$ is continuous, where the distance between ${\mathcal{F}}^{\lambda_{1}}$ and ${\mathcal{F}}^{\lambda_{2}}$ is defined as $\max\limits_{i\in{\mathcal{A}}}\left(\|f^{\lambda_{1}}_{i}-f^{\lambda_{2}}_{i}\|+\|(f^{\lambda_{1}}_{i})^{\prime}-(f^{\lambda_{2}}_{i})^{\prime}\|+\sup\limits_{x\neq y\in V}\frac{\|(f^{\lambda_{1}}_{i})^{\prime}(x)-(f^{\lambda_{2}}_{i})^{\prime}(y)\|}{|x-y|^{\theta}}\right)$. We call a parametrized family ${\mathcal{F}}^{\lambda},\lambda\in U$ a continuous family of $C^{1+\theta}$ conformal IFS. Let $\Sigma={\mathcal{A}}^{\mathbb{N}}$ be the symbolic space, and for each $\lambda\in U$, let $\Pi_{\lambda}:=\Pi_{{\mathcal{F}}^{\lambda}}:\Sigma\to\mathbb{R}^{d}$ be the corresponding natural projection map (3.2). For each $\lambda\in U$, let $\rho_{\lambda}:=\rho_{{\mathcal{F}}^{\lambda}}$ be the metric on $\Sigma$ corresponding to ${\mathcal{F}}^{\lambda}$, as defined in (3.3). Under these assumptions families $\{\Pi_{\lambda}:\lambda\in U\}$ and $\{\rho_{\lambda}:\lambda\in U\}$ satisfy assumptions (A1) and (A2). Property (A1) follows by the bounded distortion property and the distortion continuity, see for example Nakajima [Nak24, Section 3] in the more general non-autonomous case, while (A2) follows directly from the definition of metric $\rho_{\lambda}$.
@@
Let ${\mathcal{F}}=(f_{i})_{i\in{\mathcal{A}}}$ be a conformal $C^{1+\theta}$ IFS on a compact connected set $V\subset\mathbb{R}^{d}$ as described in Example 3.3 and assume that $\max\limits_{i\neq j\in{\mathcal{A}}}\|f_{i}^{\prime}\|+\|f_{j}^{\prime}\|<1$. Let $U=\{(t_{i})_{i\in{\mathcal{A}}}\in(\mathbb{R}^{d})^{{\mathcal{A}}}:f_{i}(V)+t_{i}\subset V\}$, and let $\eta$ be the normalized Lebesgue measure on $U$. Then ${\mathcal{F}}^{t}=\{f_{i}^{t}=f_{i}+t_{i}\}_{i\in{\mathcal{A}}},t\in U$ is a continuous family of $C^{1+\theta}$ conformal IFS. A family of natural projections $\{\Pi_{t}:t\in U\}$ and corresponding metrics $\{\rho_{t}:t\in U\}$ as in Example 3.7 satisfy (A1) - (A3) with measure $\eta$, with respect to the parameter $t$. See for example [BSS23, Theorem 14.5.2] for the proof in the case $d=1$, which extends in a straightforward manner to higher dimensions.
Note also that if $d=1$ and ${\mathcal{F}}_{t}=\{x\mapsto\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ consists of similarities with fixed $\lambda_{i}\in(-1,1)\setminus\{0\}$ and parametrized by $t=(t_{i})_{i\in{\mathcal{A}}}\in\mathbb{R}^{{\mathcal{A}}}$, then the similarity dimension $s_{0}=s({\mathcal{F}}_{t})$ is independent of $t$ and in fact is the unique solution of the equation $\sum\limits_{i\in{\mathcal{A}}}|\lambda_{i}|^{s_{0}}=1$. Therefore, if $s_{0}<1$, then the condition $\max\limits_{i\neq j\in{\mathcal{A}}}\|f_{i}^{\prime}\|+\|f_{j}^{\prime}\|<1$ is satisfied and hence the family ${\mathcal{F}}_{t}$ satisfies the transversality condition. Note that for $d>1$, the analogous condition $s_{0}<d$ is not sufficient for guaranteeing $\max\limits_{i\neq j\in{\mathcal{A}}}\|f_{i}^{\prime}\|+\|f_{j}^{\prime}\|<1$. ∎
Before formulating our main result on parametrized families of IFS satisfying the transversality condition, let us interpret the nearly bi-Lipschitz condition for natural projection maps as a separation condition for cylinders. Given an infinite word $\omega=(\omega_{1},\omega_{2},\ldots)\in\Sigma$, let $\omega|_{n}=(\omega_{1},\ldots,\omega_{n})$ denote its restriction to the first $n$ coordinates, while given a finite word $\omega\in\Sigma^{*}$, let $|\omega|$ denote its length and let $\omega]=\{\tau\in\Sigma:\tau|_{n}=\omega\}$ be the corresponding cylinder.
@@
Therefore, for a measure $\mu\in{\mathcal{M}}_{\mathrm{fin}}(\Sigma)$, $\mu$-a.e. $\omega\in\Sigma$ satisfies EDE if and only if $\Pi_{\mathcal{F}}$ is $\mu$-nearly bi-Lipschitz in metric $\rho_{\mathcal{F}}$. For the proof of Proposition 3.10 see Section 7.3.
Remark 3.11.
Let us emphasize that EDE is a much weaker separation condition that the Open Set Condition (see [BSS23, Definition 1.5.2]), so that it can hold for typical IFS with overlaps, while it is strong enough so that it can be used to deduce novel results on the multifractal analysis, see Section 8. In particular, see Lemma 8.1 for a useful consequence of EDE, which is the main ingredient of the proof of Theorem 1.1.
The following is our main result on transversal families of conformal IFS.
@@
$\begin{split}&\gamma\times\eta\left(\left\{(V,\lambda):\ \|P_{V}\Pi_{\lambda}(\omega)-P_{V}\Pi_{\lambda}(\tau)\|<\rho_{\lambda}(\omega,\tau)r\right\}\right)\\ &\quad=\int\gamma\left(\left\{V:\|P_{V}\Pi_{\lambda}(\omega)-P_{V}\Pi_{\lambda}(\tau)\|<\|\Pi_{\lambda}(\omega)-\Pi_{\lambda}(\tau)\|\frac{\rho_{\lambda}(\omega,\tau)r}{\|\Pi_{\lambda}(\omega)-\Pi_{\lambda}(\tau)\|}\right\}\right)d\eta(\lambda)\\ &\quad\leq c\int\left(\frac{\rho_{\lambda}(\omega,\tau)r}{\|\Pi_{\lambda}(\omega)-\Pi_{\lambda}(\tau)\|}\right)^{d}d\eta(\lambda)=cr^{d}\int_{0}^{\infty}\eta\left(\left\{\lambda:\left(\frac{\rho_{\lambda}(\omega,\tau)r}{\|\Pi_{\lambda}(\omega)-\Pi_{\lambda}(\tau)\|}\right)^{d}>a\right\}\right)da\\ &\quad\leq c^{\prime}r^{d}\int_{0}^{\infty}\min\left\{1,a^{-n/d}\right\}da\leq c^{\prime\prime}r^{d}.\end{split}$